In a population of 50,000 diploid individuals, what is the probability that a new neutral mutation will ultimately reach fixation?
  
   Yes, that's correct!
  
  
a.     0.001
  
b.     0.0001
  
c.     0.00001
  
d.     0.000001
  
e.     0.0000001
Answer #1

Probability of fixation of neutral mutation = 1/2N

N is 50,000

Probability = 1/ (2×50,000) = 1 / 100,000 = 0.00001

Option c is correct.

As per the theory of molecular evolution, probability of fixation of an allele depends upon its,

Initial frequency

Selective advantage or disadvantage

Effective size of population

Number of mutants arising for a given population is given by 2N.

Please rate.